OpenStudy (anonymous):

The product of two integers is -21. The difference between the integers is -10. The sum of the two integers is 4. What are the two integers?

OpenStudy (midhun.madhu1987):

The product of 2 integers = -21 Let one integer be x and the other will be -21/x Difference = -10 , so x - (-21/x) = -10 Sum = 4 , so x + (-21/x) = 4 Find x using these quations...

OpenStudy (anonymous):

{x y = -21, x - y = -10, x + y = 4} {x = -3, y = 7}